OpenStudy (anonymous):

HHHEEEEELLLLLPPPPPPPPPPPPP!!! (1+cot^2)/sec^2

OpenStudy (anonymous):

\[\frac{ 1+\cot^2\theta }{ \sec^2\theta }\]